OpenStudy (anonymous):

Evaluate the following limits algebraically: lim x -> infinity 7x^3 + 5x/(8x^9 + x^6 + 4) ^1/3

OpenStudy (kirbykirby):

\[ \lim_{x\to\infty}7x^3+\frac{5x}{(8x^9+x^6+4)^{1/3}}\] ?

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} \frac{ 7x ^{3} + 5x}{ (8x^9 + x^6 + 4) ^1/3}\]

OpenStudy (kirbykirby):

Oh ok \[ \large \begin{align}\lim_{x\to\infty}\frac{7x^3+5x}{(8x^9+x^6+4)^{1/3}}&=\lim_{x\to\infty}\frac{x^3\left( 7+\frac{5}{x^2}\right)}{\left(x^9\left(8+\frac{1}{x^3}+\frac{4}{x^9} \right) \right)^{1/3}}\\&=\lim_{x\to\infty}\frac{x^3\left( 7+\frac{5}{x^2}\right)}{\left( 8+\frac{1}{x^3}+\frac{4}{x^9}\right)^{1/3}x^3}\\&=\lim_{x\to\infty}\frac{\left( 7+\frac{5}{x^2}\right)}{\left( 8+\frac{1}{x^3}+\frac{4}{x^9}\right)^{1/3}} \\&=\frac{7}{8^{1/3}}=\frac{7}{2} \end{align}\] The before last line simplifies because all the limits for fractions with x in the denominator go to 0 as x goes to infinity

OpenStudy (mathmath333):

\(\Large\tt \color{black}{great~~~~\ddot\smile}\)

OpenStudy (kirbykirby):

\(\dddot \smile\) alien smile

OpenStudy (mathmath333):

lol

OpenStudy (anonymous):

thank you both so much!!! :)

OpenStudy (mathmath333):

yw\(\large\tt \color{black}{great\ddddot\smile}\)

OpenStudy (kirbykirby):

=]